A rop of length l slides with it’s ends on co-ordinates axis. Find the locus of it’s midpoint.

Grade:12

1 Answers

RITABRATA SANYAL
9 years ago
see, if (a,b)are the x n y axes,a2+b2=l2.....(c,d) midpt....c=a/2,,d=a/2....c^2+d^2=l^2/4....x^2+y^2=l2/4